You must evaluate the purchase of a proposed spectrometer for the R&D department. The purchase price of the spectrometer including modifications is $160,000, and the equipment will be fully depreciated at the time of purchase. The equipment would be sold after 3 years for $61,000. The equipment would require a $6,000 increase in net operating working capital (spare parts inventory). The project would have no effect on revenues, but it should save the firm $51,000 per year in before-tax labor costs. The firm's marginal federal-plus-state tax rate is 25%.

  1. What is the initial investment outlay for the spectrometer, that is, what is the Year 0 project cash flow? Enter your answer as a positive value. Round your answer to the nearest dollar.
    $  

  2. What are the project's annual cash flows in Years 1, 2, and 3? Do not round intermediate calculations. Round your answers to the nearest dollar.
    Year 1: $   
    Year 2: $   
    Year 3: $  

  3. If the WACC is 14%, should the spectrometer be purchased?
    -Select-Yes No
Answer #1

a)

Calculation of Initial investment:

Initial investment on machine = purchase price = $160,000

Initial working capital required = $6,000

Total Initial Investment = $ 160,000 + $6,000 = $166,000

c)

Asset landed price = purchase price = $160,000

As the asset is fully depreciated at the time of purchase itself, no further depreciation is calculated.

Proceeds/profits from Asset Sale = purchase price – salvage value

Salvage value = $61,000

Profit from asset sale = $160,000- $61,000 = $99,000

Tax applied = $99,000*25% = $24,750

Net income from asset sale = $61,000- $24,750= $36,250

Calculation for Project Cash Flow:

Year 0 cash flow = Initial outlay = -$298,000

Year 1 cash flow = income in terms of after tax cost savings + tax savings in terms of depreciation

= $51,000 * (1-25%)+ 0 = $38,250

Similarly, Year 2 cash flow = $51,000 * (1-25%)+ 0 = $38,250

Year 3 cash flow = income in terms of after tax cost savings + tax saving in terms of depreciation + net income from asset sale + recovery of working capital

$51,000 * (1-25%)+ 0 + $36,250+ $6,000 = $38,250 + 0 + $36,250+ $6,000 = $80,500

e)

In order to decide whether to buy or not, we need to calculate NPV of the project

NPV @ WACC 14% = -$ 166,000 + $38,250/(1.14) + $38,250 / (1.14^2) + $80,500 / (1.14^3)

NPV = -$ 166,000 + $33,552.6316 + $29,432.133 + $54,335.2071 = -$ 166,000 + $117,319.9717

NPV = -$48,680.0283

As NPV<0, the machine should not to be purchased.

So the answer is NO.